5/22-23: MZ release from PSL
5/20-24: Exchanging cryo valves.
5/27-6/2: Y arm mass lock. (not done yet)
5/29: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (specially BS)
- 6/3-6/14: Installation of TMS for ETMY
- 6/8: Engineering run
5/22AM~: tuned off cryostat cooler, finesse measurement for Xarm
5/22AM~: tuned off shield cooler, finesse measurement for Xarm
5/2: Cooling start
5/23AM~: tuned off cryostat cooler, finesse measurement for Yarm
5/17~: tuned off ETMY cryostat cooler
- Long term:
- 4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- TMS installation this week.
- Please reduce the number or speed of 16kHz channels.
Report of work
- COM: PRMI robust lock was established, with free swinging SRM.
- COM: Very short DRMI lock was established. Glitch of somewhere was seen in PRCL and SRCL.
- COM: SRY was locked.
- Type A: ITMY is still shaky. Investigation is planned.
- AOS: TMSY was installed and aligned with transmittance of GR.
- It will be done with light with Y arm lock next.
- Found LVDT issue.
- CAL: Reconstruction monitor for X arm was done with calibration of MICH. OLG and other TF ware measured.
- COM: Centering oplev, BS oplev investigation, alignment of AS.
- IOO, FCL: Filters will be stopped at 2PM around IOO for tomorrows run. No stop around PRM or SRM and ETXM.